OpenStudy (anonymous):

Solve for x. x-9=3x+11

OpenStudy (amistre64):

what methods do you know of that will help you in this?

OpenStudy (anonymous):

subtract 3x from both sides, then add 9 to both sides, finally diivde -2

OpenStudy (amistre64):

the process is more important than the answer at the moment; these problem are designed to give you the practice you need to develop good math skills later on

OpenStudy (anonymous):

x=3x+11+9 x-3x=20 -2x=20 x=20/-2 x=-10

OpenStudy (anonymous):

You are right I did mine way too fast!

OpenStudy (anonymous):

x=3x+11+9 x-3x=20 -2x=20 x=20/-2 x=-10

OpenStudy (amistre64):

even when i do mine slow it goes awry some how ;)